Contact: Rita Baysinger, Public Information Officer, (303) 345-8576, firstname.lastname@example.org
Date Started: June 20, 2105
Current Acreage: 1,800 no change
Management Option: Full
July 5: Fuels continued to dry. The fire activity was moderate today with single tree torching. The crews will continue to prep the dozer line for a possible burnout that is scheduled for mid-week. Fire activity yesterday was moderate with backing and creeping. There was some smoke visible in and near Tok. Fire received no moisture with daytime temps in the 80s.
July 6: With highs predicted to be in the 80s today and gusts to 12 expect fire behavior to be more active today with potential spread to the North. Objectives continue to be to minimize impacts on tribal lands and to prevent spread in the direction of Tok. 106 personnel on the fire yesterday.